Shift summary DAY 01/22/00


Subject: Shift summary DAY 01/22/00
From: Vivian O'Dell (odell@fnal.gov)
Date: Fri Feb 04 2000 - 15:36:13 CST


             Shift Summary for day shift, 01/22/00

On Shift: R. Ford

Kami cooldown in progress. Lots of work completed in NM2.

        0730 Lots of workers show up for work in NM2.
                Walkthru complete

        10:00 TRD gas_mon is down, restart.

        1100 Completed walk thru. TRD gas_mon showing CAMAC errors. Paged
                Sasha G. He found crates 10 & 13 turned off which needed to
                be on for TRD gas_mon to work. Also did SDAQ fresh_start.

        1456 Completed walk thru. NM2 work going well

        1620 Crews finished raising NM2 magnets. Survey work continues.
                Tomorrow vacuum work will be done. Electricians start on
                Mon.

========================================================================
Follow up Notes on Previous Shift
        FB off.

========================================================================
Items to Follow up on Next Shift
        Relax, not much is going to happen.



This archive was generated by hypermail 2b27 : Fri Feb 04 2000 - 15:36:15 CST